OpenStudy (anonymous):

If x=b+4 & y=b-3 then in terms of x and y, b = show steps thanks

OpenStudy (anonymous):

I think you need another equations to do this because there are two equations and three unknowns. b/c i get x-4=b y+3=b

OpenStudy (gw2011):

I agree with the above answer. Here are the steps: x=b+4 x-4=b+4-4 x-4=b y=b-3 y+3=b-3+3 y+3=b